# The iterated Dirichlet process and applications to Bayesian inference

### v1

@@ Abstract
Consider an i.i.d. sequence of random variables, taking values in some space $S$, whose underlying distribution is unknown. In problems of Bayesian inference, one models this unknown distribution as a random measure, and the law of this random measure is the prior. When $S=\{0,1\}$, a commonly used prior is the uniform distribution on $[0,1]$, or more generally, the beta distribution. When $S$ is finite, the analogous choice is the Dirichlet distribution. For a general space $S$, we are led naturally to the Dirichlet process (see Ferguson [5]).
Here, we consider an array of random variables, and in so doing are led to what we call the iterated Dirichlet process (IDP). We define the IDP and then show how to compute the posterior distribution, given a finite set of observations, using the method of sequential imputation. Ordinarily, this method requires the existence of certain joint density functions, which the IDP lacks. We therefore present a new, more general proof of the validity of sequential imputation, and show that the hypotheses of our proof are satisfied by the IDP.
@@ 1 Introduction
Let $S$ be a complete and separable metric space and $\{\eta_{i}:i\in\mathbb{N}\}$ an exchangeable sequence of $S$-valued random variables. By de Finetti’s theorem, there exists a random measure $\lambda$ on $S$ such that, given $\lambda$, the sequence $\eta_{1},\eta_{2},\ldots$ is conditionally i.i.d. with distribution $\lambda$. To compute the conditional distribution of $\eta_{N+1},\eta_{N+2},\ldots$ given $\eta_{1},\ldots,\eta_{N}$, we must, among other things, assign a prior distribution to $\lambda$.
The random measure $\lambda$ is an element of $M_{1}=M_{1}(S)$, the space of probability measures on $S$. A prior distribution for $\lambda$ is therefore an element of $M_{1}(M_{1}(S))$. In the case $S=\{0,1\}$, we can identify a measure $\nu\in M_{1}(S)$ with its projected value $\nu(\{1\})\in[0,1]$. In this way, we can identify $M_{1}(M_{1}(S))$ with $M_{1}([0,1])$. Hence, a prior distribution for $\lambda$ is simply a probability measure on $[0,1]$. A common choice for a non-informative prior in this case is the uniform distribution, or its generalization, the beta distribution.
In the case $S=\{0,\ldots,d\}$, we can identify a measure $\nu\in M_{1}(S)$ with its projected values $(\nu(\{0\}),\ldots,\nu(\{d\}))\in\Delta^{d}$, where $\Delta^{d}$ is the standard $d$-simplex,
${\Delta^{d}=\{t=(t_{0},\ldots,t_{d})\in\mathbb{R}^{d+1}:t_{j}\geq 0\text{ and }\sum_{0}^{d}t_{j}=1\}.}$ (1.1)
In this way, a prior distribution for $\lambda$ is simply a probability measure on $\Delta^{d}$. A common choice in this case is the Dirichlet distribution.
In the case of a general $S$, a common choice for a non-informative, non-parametric prior is to let $\lambda$ be a Dirichlet process on $S$. The Dirichlet process was first introduced by Thomas S. Ferguson in [5]. Its definition and relevant properties are formally summarized in Section 3. A Dirichlet process is almost surely a discrete measure. It has two parameters, $\kappa$ and $\rho$. The probability measure $\rho\in M_{1}(S)$ is its mean measure, or base measure, satisfying $E[\lambda(A)]=\rho(A)$. The number $\kappa\in(0,\infty)$ is its concentration parameter. The smaller $\kappa$ is, the more likely $\lambda$ is to be concentrated on only a few points in $S$. If $\lambda$ is a Dirichlet process on $S$ with parameters $\kappa$ and $\rho$, then its law, which is an element of $M_{1}(M_{1}(S))$, is denoted by $\mathcal{D}(\kappa\rho)$.
Now let $\{\xi_{ij}:i,j\in\mathbb{N}\}$ be an array of $S$-valued random variables. In [4], such arrays were intended to model a situation in which there are several agents, all from the same population, each undertaking a sequence of actions. The space $S$ represents the set of possible actions and $\xi_{ij}$ represents the $j$th action of the $i$th agent. A simple example presented in [4] was the following. Imagine a pressed penny machine, like those found in museums or tourist attractions. For a fee, the machine presses a penny into a commemorative souvenir. Now imagine the machine is broken, so that it mangles all the pennies we feed it. Each pressed penny it creates is mangled in its own way, and has its own probability of landing on heads when flipped. The machine, though, might have certain tendencies. For instance, it might tend to produce pennies that are biased toward heads. In this situation, the agents are the pennies and the actions are the heads and tails that they produce. In this example, then, we have $S=\{0,1\}$ and the random variable $\xi_{ij}\in S$ denotes the result of the $j$th flip of the $i$th penny created by the machine.
We assume that $\xi$ is row exchangeable, which means that
@@
    each row of the array, $\xi_{i}=\{\xi_{ij}:j\in\mathbb{N}\}$, is an exchangeable sequence of $S$-valued random variables, and
@@
    the sequence of rows, $\{\xi_{i}:i\in\mathbb{N}\}$, is an exchangeable sequence of $S^{\infty}$-valued random variables.
It is easy to see that $\xi$ is row exchangeable if and only if $\{\xi_{\sigma(i),\tau_{i}(j)}\}$ and $\{\xi_{ij}\}$ have the same finite-dimensional distributions whenever $\sigma$ and $\tau_{i}$ are (finite) permutations of $\mathbb{N}$. As a matter of notation, we will sometimes write $\xi(i)$ and $\xi(i,j)$ for $\xi_{i}$ and $\xi_{ij}$, respectively.
According to a de Finetti theorem for row exchangeable arrays (see [4, Theorem 3.2]), there exists a sequence of random measure $\mu_{1},\mu_{2},\ldots$ on $S$ and a random measure $\varpi$ on $M_{1}(S)$ such that
@@
The random measures $\mu_{i}$ are called the row distributions of $\xi$, and $\varpi$ is called the row distribution generator.
In the present work, we are concerned with performing Bayesian inference on row exchangeable arrays. The first step toward this to assign a prior distribution to $\xi$. According to the above, this is equivalent to assigning to a prior distribution to $\varpi$. Following Ferguson, we would like to define $\varpi$ to be a Dirichlet process. Since $\varpi$ is a random measure on $M_{1}(S)$, the law of $\varpi$ should have the form $\mathcal{D}(\kappa\rho)$, where $\kappa>0$ and $\rho$ is a nonrandom measure on $M_{1}(S)$. If $B\subseteq M_{1}(S)$ is measurable, then
@@
In other words, $\rho$ is the prior distribution for $\mu_{i}$, which is a random measure on $S$. Again turning to the Dirichlet process, we let $\rho=\mathcal{D}(\varepsilon\varrho)$, where $\varepsilon>0$ and $\varrho$ is a nonrandom measure on $S$. As above, if $A\subseteq S$ is measurable, then
$P(\xi_{ij}\in A)=E[P(\xi_{ij}\in A\mid\mu_{i})]=E[\mu_{i}(A)]=\varrho(A),$
so that $\varrho$ is the prior distribution for $\xi_{ij}$. This means that $\varpi$ has distribution $\mathcal{D}(\kappa\mathcal{D}(\varepsilon\varrho))$. We call a random measure with a distribution of this form an iterated Dirichlet process (or IDP) on $S$ with parameters $\kappa$, $\varepsilon$, and $\varrho$. We call $\kappa$ and $\varepsilon$ the column and row concentrations, respectively, and we call $\varrho$ the base measure of $\varpi$. The effect of $\kappa$ and $\varepsilon$ on the behavior of $\varpi$ is discussed in Section 3.5.
Having assigned a prior, we turn our attention to inference. We adopt the notation used in [4]:
@@
The exact computation of (1.2) is intractable. For example, consider the simplest nontrivial case, $M=2$ and $S=\{0,1\}$. When $S=\{0,1\}$, an IDP on $S$ can be regarded as simply a Dirichlet process on $[0,1]$ whose base measure is a Beta distribution. Such a process is considered in [3]. There, the authors provide us with a formula for
$P(\xi_{2,N+1}=1\mid\xi_{ij}=k_{ij},i=1,2,j=1,\ldots,N).$
To state this formula, let $\varrho_{k}=\varrho(\{k\})$, $\alpha_{k}=\varepsilon\varrho_{k}$, and $y_{i}=\sum_{j=1}^{N}k_{ij}$. Using the notation $t^{(n)}=\Gamma(t+n)/\Gamma(t)=\prod_{k=0}^{n-1}(t+k)$, define
$r={\frac{\kappa\varepsilon^{(2N)}}{(\varepsilon^{(N)})^{2}}\frac{\alpha_{1}^{(y_{1})}\alpha_{1}^{(y_{2})}}{\alpha_{1}^{(y_{1}+y_{2})}}\frac{\alpha_{0}^{(N-y_{1})}\alpha_{0}^{(N-y_{2})}}{\alpha_{0}^{(2N-y_{1}-y_{2})}}}.$
We then have
$P(\xi_{2,N+1}=1\mid\xi_{ij}=k_{ij},i=1,2,j=1,\ldots,N)\\ ={\frac{r}{1+r}\left(\frac{\varepsilon}{\varepsilon+N}\varrho_{1}+\frac{N}{\varepsilon+n}\frac{y_{2}}{N}\right)+\frac{1}{1+r}\left(\frac{\varepsilon}{\varepsilon+2N}\varrho_{1}+\frac{2N}{\varepsilon+2N}\frac{y_{1}+y_{2}}{2N}\right)}.$
The case $M=3$ and $S=\{0,1\}$ is also outlined in [3], and involves roughly twice as many computations. To deal with larger values of $M$, as well as more general spaces $S$, we need to adopt a different approach to the computation of (1.2).
In [4], we showed that (1.2) can be computed in terms of the distributions,
@@
where $1\leq m\leq M$ (see Section 2.4). But the same issue that makes (1.2) intractable also hinders the computation of (1.3). Namely, in both cases, we are trying to condition on more than one row of the array $\xi$. If, instead of (1.3), we wanted to compute
@@
for $1\leq m\leq M$, then this can be done (see Section 5.3). The problem, then, is to find a way to use (1.4) to determine (1.2).
This was done by Liu in [12] for the special case $S=\{0,1\}$, using a method known as sequential imputation. The proofs in [12] cite [11], in which Kong, Liu, and Wong first developed the method of sequential imputation. In [11, Section 2], the authors give their core argument, demonstrating the validity of the method. This argument evidently presupposes that the underlying random variables have densities with respect to some product measure. In Proposition 4.5, we show that this is not true for the IDP, not even in the case $S=\{0,1\}$ that was treated by Liu in [12]. In other words, the work in [11] is not sufficient to justify its use in [12].
In Theorem 4.7, we provide a new proof that justifies sequential imputation. Our proof avoids the supposition of a full joint density. In doing so, it shows that sequential imputation can be validly applied to an IDP on any complete and separable metric space. This includes the special case $S=\{0,1\}$ treated in [12] as well as the general case treated here.
When we apply sequential imputation to the IDP, we arrive at our main result, Theorem 5.2. Combined with [4, Theorem 3.4], which is presented below as Theorem 2.4, this allows us to compute (1.2).
The outline of this paper is as follows. In Section 2, we present some basic background results and establish notation. We also include the main results from [4] on row exchangeable arrays. In Section 3, we give the necessary background on Dirichlet processes, including Ferguson’s definition, the Sethuraman stick-breaking construction, mixtures of Dirichlet processes, and the formulas of inference for both ordinary observations as well as noisy observations.
In Section 3.5, we take a closer look at the iterated Dirichlet process, which we defined above. Namely, we discuss the relationship between the parameters, $\kappa$, $\varepsilon$, and $\varrho$, and the behavior of $\varpi$. It should be noted that the iterated Dirichlet process is not the same thing as the hierarchical Dirichlet process, a process that appears elsewhere in the literature (see, for example, [15, 16]). The latter is a mixture of Dirichlet processes, whereas the IDP is not. For a more detailed explanation of the differences between the two, see Remark 3.6.
In Section 4, as described above, we discuss the method of sequential imputation. We then apply this in Section 5 to the iterated Dirichlet process, obtaining our main results, which are Theorem 5.2 and Corollary 5.3. Finally, in Section 6, we present several hypothetical examples to illustrate the use of our main results.
2 Background and notation
2.1 Random measures and kernels
If $(X,\mathcal{U})$ is a topological space, then $\mathcal{B}(X)=\sigma(\mathcal{U})$ denotes the Borel $\sigma$-algebra on $X$. We may sometimes denote $\mathcal{B}(X)$ by $\mathcal{B}_{X}$. We use $\mathcal{R}^{d}$ to denote the Borel $\sigma$-algebra on $\mathbb{R}^{d}$ with the Euclidean topology. If $(S,\mathcal{S})$ is a measurable space and $B\in\mathcal{S}$, then $\mathcal{S}|_{B}=\{A\cap B:A\in\mathcal{S}\}=\{A\in\mathcal{S}:A\subseteq B\}$. We let $\mathbb{R}^{*}=[-\infty,\infty]$ denote the extended real line, equipped with the topology generated by the metric $\rho(x,y)=|\tan^{-1}(x)-\tan^{-1}(y)|$. We use $\mathcal{R}^{*}$ to denote $\mathcal{B}(\mathbb{R}^{*})$. Note that =$\mathcal{R}^{*}=\{A\subseteq\mathbb{R}^{*}:A\cap\mathbb{R}\in\mathcal{R}\}$.
[…]
Let $M_{1}=M_{1}(S)$ be the set of all probability measures on $S$. Note that $M_{1}=\{\nu\in M:\nu(S)\in\{1\}\}$. Hence, $M_{1}\in\mathcal{M}$ and we may define $\mathcal{M}_{1}$ by $\mathcal{M}_{1}=\mathcal{M}|_{M_{1}}$. A random probability measure on $S$ is a function $\mu:\Omega\to M_{1}$ that is $(\mathcal{F},\mathcal{M}_{1})$-measurable, or equivalently, a random measure taking values in $M_{1}$.
We equip $M_{1}$ with the Prohorov metric, $\bm{\pi}$, which metrizes weak convergence. Since $S$ is complete and separable, $M_{1}$ is complete and separable under $\bm{\pi}$. It can be shown that $\mathcal{M}_{1}=\mathcal{B}(M_{1})$. (See, for example, [6, Theorem 2.3].)
Unless otherwise specified, we will equip $S^{\infty}$ with the metric ${\bf d}^{\infty}$ defined by
${\bf d}^{\infty}(x,y)=\sum_{n=1}^{\infty}\frac{{\bf d}(x_{n},y_{n})\wedge 1}{2^{n}}.$
Note that $x\to y$ in $S^{\infty}$ if and only if $x_{n}\to y_{n}$ in $S$ for all $n$. In particular, the metric ${\bf d}^{\infty}$ induces the product topology on $S^{\infty}$. Consequently, $\mathcal{S}^{\infty}=\mathcal{B}(S^{\infty})$ and $S^{\infty}$ is separable. In fact, $(S^{\infty},{\bf d}^{\infty})$ is a complete and separable metric space.
Let $(T,\mathcal{T})$ be a measurable space. A function $\mu:T\to M$ is called a kernel from $T$ to $S$ if $\mu(\cdot,B)$ is $(\mathcal{T},\mathcal{R}^{*})$-measurable for all $B\in\mathcal{S}$. By [7, Lemma 1.37], given a function $\mu:T\to M$ and a $\pi$-system $\mathcal{C}$ such that $\mathcal{S}=\sigma(\mathcal{C})$, the following are equivalent:
  1. (i)
    $\mu$ is a kernel.
  2. (ii)
    $\mu(\cdot,B)$ is $(\mathcal{T},\mathcal{R}^{*})$-measurable for all $B\in\mathcal{C}$.

### v2

# Conditional distributions for the nested Dirichlet process via sequential imputation

@@ Abstract
We consider an array of random variables, taking values in a complete and separable metric space, that exhibits a kind of symmetry which we call row exchangeability. Given such an array, a natural model for Bayesian nonparametric inference is the nested Dirichlet process (NDP). Exactly determining posterior distributions for the NDP is infeasible, since the computations involved grow exponentially with the sample size. In this paper, we present a new approach to determining these posterior distributions that involves the use of sequential imputation.
@@ 1 Introduction
1.1 Motivation and main objective
Consider a situation in which there are several agents, all from the same population. Each agent undertakes a sequence of actions. These actions are chosen according to the agent’s particular tendencies. Although different agents have different tendencies, there may be patterns in the population.
  • The future behavior of the agents we have observed.
  • The behavior of a new (unobserved) agent from the population.
Let $S$ denote the set of possible actions the agents may take. We assume that $S$ is a complete and separable metric space. Let $\xi_{ij}$ denote the $j$th action by the $i$th agent. We assume that $\{\xi_{\sigma(i),\tau_{i}(j)}\}$ and $\{\xi_{ij}\}$ have the same finite-dimensional distributions whenever $\sigma$ and $\tau_{i}$ are permutations. We will say that such an array is row exchangeable.
Note that if $\xi=\{\xi_{ij}\}$ is row exchangeable, then for each $i$, the sequence $\xi_{i}=\{\xi_{ij}:j\in\mathbb{N}\}$ is an exchangeable sequence of $S$-valued random variables, and that the sequence of sequences, $\xi=\{\xi_{i}:i\in\mathbb{N}\}$ is also exchangeable. Let $M_{1}(S)$ denote the set of probability measures on $S$. We equip $M_{1}(S)$ with the Prohorov metric, so that $M_{1}(S)$ is also a complete and separable metric space. By de Finetti’s theorem, there exists a sequence of random Borel probability measures $\mu_{1},\mu_{2},\ldots$ on $S$ and a random Borel probability measure $\varpi$ on $M_{1}(S)$ such that
@@ 1.1 Motivation and main objective
We call the $\mu_{i}$ the row distributions of the random array $\xi=\{\xi_{ij}\}$, and we call $\varpi$ the row distribution generator.
Our goal is to make inferences about the array $\xi$ based on observations of some of its entries. We wish for this inference to be both Bayesian and nonparametric. To facilitate Bayesian inference, we must place a prior distribution on the random measure $\varpi$. We make the nonparametric choice of letting $\varpi$ be a Dirichlet process. That is, $\varpi\sim\mathcal{D}(\kappa\rho)$ for some $\kappa>0$ and some Borel probability measure $\rho$ on $M_{1}(S)$. To choose the measure $\rho$, we first observe that
@@
Hence, the measure $\rho$ is the prior distribution for $\mu_{i}$. In keeping with our aim of nonparametric inference, we also let $\rho$ be the distribution of a Dirichlet process. That is, $\rho=\mathcal{D}(\varepsilon\varrho)$ for some $\varepsilon>0$ and some Borel probability measure $\varrho$ on $S$.
This gives us the model $\varpi\sim\mathcal{D}(\kappa\mathcal{D}(\varepsilon\varrho))$. In other words, the process $\varpi$ is what Rodríguez, Dunson, and Gelfand call a nested Dirichlet process (NDP) in [17]. In that paper, the authors use a motivating example in which the agents are different medical centers, and the actions are the individual patient outcomes produced by these centers.
We call $\kappa$ and $\varepsilon$ the column and row concentrations of $\varpi$, respectively, and we call $\varrho$ the base measure of $\varpi$. We adopt the following notation:
@@
As demonstrated in [17], algorithms based on Pólya urns are infeasible in this situation. They require evaluating distributions where the number of terms grows exponentially with the sample size. For the same reason, the exact computation of (1.1) is infeasible. (The exact computation in the simplest nontrivial case, $M=2$ and $S=\{0,1\}$, is given in [4] and takes up a full page.) In [17], the authors deal with the infeasibility of posterior computation by using truncation. Here, we take a different approach, motivated by the work of Liu and coauthors in [10, 11]. In [11], Liu considered what is effectively an NDP with $S=\{0,1\}$. Using the method of sequential imputation, developed in [10], Liu was able to determine the posterior distributions (1.1) without truncation. Unfortunately, Liu’s proof contains a fatal flaw. In this paper, we correct that flaw, then generalize the method to arbitrary $S$.
1.2 The role of sequential imputation
To explain how sequential imputation enters into the determination of (1.1), consider the following. It is straightforward to verify that
$P\left(\bigcap_{i=1}^{M}\bigcap_{j=N_{i}+1}^{O_{i}}\{\xi_{ij}\in A_{ij}\}\;\middle|\;\mathcal{G}\right)=E\left[\prod_{i=1}^{M}\prod_{j=N_{i}+1}^{O_{i}}\mu_{i}(A_{ij})\;\middle|\;\mathcal{G}\right],$
for all Borel sets $A_{ij}\subseteq S$ and any sub-$\sigma$-algebra $\mathcal{G}\subseteq\sigma(X_{1N_{1}},X_{2N_{2}},\ldots,X_{MN_{M}})$. Taking $N_{i}=N$, $O_{i}=N^{\prime}$, and $\mathcal{G}=\sigma({\bm{X}}_{MN})$, we see that (1.1) is entirely determined by $\mathcal{L}(\bm{\mu}_{M}\mid{\bm{X}}_{MN})$. Note that
In general, ${\bm{X}}_{mN}$ and $\{(\mu_{j},X_{jN})\}_{j=m+1}^{M}$ are conditionally independent given $\bm{\mu}_{m}$. Hence, in (1.2), the first m1m-1 rows of ${\bm{X}}_{MN}$ can be omitted. In other words, the posterior distribution (1.1) is entirely determined by the conditional distributions,
@@ 1.2 The role of sequential imputation
where $1\leq m\leq M$.
When $m<M$, the determination of (1.3) involves conditioning on more than one row of the array $\xi$. This is exactly the issue that makes Pólya-urn-based algorithms and exact computations of (1.1) infeasible. The number of required calculations grows exponentially with the number of rows. If, instead of (1.3), we wanted to compute
@@
for $1\leq m\leq M$, then this is feasible. The problem, then, is to find a way to use (1.4) to determine (1.1).
This is where sequential imputation is used. The entire row of data, $\xi_{m}=\{\xi_{mj}\}_{j=1}^{\infty}$, is enough to determine $\mu_{m}$. But in (1.4), we observe $X_{mN}$, which is only part of this row of data. Hence, we are faced with a sequence, indexed by $m$, of missing data problems. This is precisely the situation that sequential imputation is designed to handle.
1.3 Other related models
1.3.1 The dependent Dirichlet process
Although the NDP arises naturally when considering row exchangeable arrays, it is just one of many models used for nonparametric Bayesian inference. In the NDP, we see that we have a sequence $\mu=\{\mu_{i}\}_{i=1}^{\infty}$ of dependent Dirichlet processes. The study of dependent Dirichlet processes goes back at least to [12, 13]. Since a Dirichlet process is almost surely a discrete measure, it is characterized by its atoms (the countable set of points on which it is supported) and its weights (the amount of mass it assigns to each atom). In [12, 13], a sequence of dependent Dirichlet processes is defined by a specific construction of the weights and atoms. The resulting process has come to be known by the name, dependent Dirichlet process (DDP). It should be noted, though, that not every family of dependent Dirichlet processes is a DDP. Our sequence $\mu$, for instance, is not a DDP, but is rather a variation of the DDP. In [2], an alternative, equivalent definition of the DDP was given in terms of copulas. For a survey of the DDP and related models, see [16].
Two common categories of DDPs are the single-weights DDPs and the single-atoms DDPs. In the former, all the Dirichlet processes in the DDP share the same weights; in the latter, they share the same atoms. The NDP does not fit into either of these categories. In fact, as long as $\varrho$ is non-atomic, either $\mu_{i}=\mu_{j}$ or $\mu_{i}$ and $\mu_{j}$ have no atoms and no weights in common.
1.3.2 Mixture models
There are a number of popular variations of the DDP. For instance, in [5], the authors propose a model which begins with a family of independent Dirichlet processes. The dependent Dirichlet processes are then constructed as weighted mixtures of the independent ones, with the dependence structure determined by the weights. In a slightly different approach, [14] proposes a model in which the dependence is created by a common underlying Dirichlet process. That is, $\mu_{j}=c\widetilde{\mu}_{0}+(1-c)\widetilde{\mu}_{j}$, where $\{\widetilde{\mu}_{j}\}_{j=0}^{\infty}$ are independent Dirichlet processes. The parameter $c$ allows for control over the degree of dependence among the different $\mu_{j}$. The authors call this a hierarchical Dirichlet process mixture model.
The hierarchical Dirichlet process mixture of [14] should not be confused with the hierarchical Dirichlet process (HDP), which was introduced in [19]. Compared to the mixture processes of [5] and [14], the HDP seems, at least on the surface, to be much more closely related to the NDP. In fact, though, they are quite different. The HDP is a Dirichlet process whose base measure is a Dirichlet process. The NDP, however, is a Dirichlet process whose base measure is the law of a Dirichlet process. If we are not careful about distinguishing between a process and its law, then we could easily mistake one for the other.
The HDP and NDP, in fact, have different state spaces. The HDP takes values in $M_{1}(S)$, whereas the NDP takes values in $M_{1}(M_{1}(S))$. For example, if we take $\kappa^{\prime}$ to be random and $\rho^{\prime}\sim\mathcal{D}(\varepsilon\varrho)$, then we can define $\lambda$ to be an HDP by using the conditional distribution ${\lambda\mid\kappa^{\prime},\rho^{\prime}}\sim\mathcal{D}(\kappa^{\prime}\rho^{\prime})$. Note that $\rho^{\prime}$ is an $M_{1}(S)$-valued random variable. In contrast, for the NDP, we take $\kappa$ and $\rho$ to be nonrandom and $\rho=\mathcal{D}(\varepsilon\varrho)$. In this case, $\rho$ is a nonrandom element of $M_{1}(M_{1}(S))$. We then define $\varpi$ to be an NDP by the unconditional distribution $\varpi\sim\mathcal{D}(\kappa\rho)$.
1.3.3 The infinite relational model
A model that does bear a close connection to the NDP is the infinite relational model (IRM), introduced independently in both [7] and [20]. It is a cluster-based model that can be regarded as a kind of nonparametric stochastic block model, and is common in the machine learning literature. For a survey of the IRM and other Bayesian models of exchangeable structures, see [15].
We could obtain an IRM from our process $\xi$ by only a slight modification. To understand this modification, we should clarify that although (i) and (ii) above are consequences of the row exchangeability of $\xi$, they do not characterize row exchangeability. For example, an array $\xi$ is said to be separately exchangeable if $\{\xi_{\sigma(i),\tau(j)}\}$ and $\{\xi_{ij}\}$ have the same finite-dimensional distributions whenever $\sigma$ and $\tau$ are permutations. A separately exchangeable array also satisfies (i) and (ii). But a row exchangeable array satisfies